OpenStudy (anonymous):

If the diameter of a student’s brain is about 6 inches, and you assume its shape is approximately a hemisphere, then what is the volume of the student’s brain?

OpenStudy (anonymous):

do you know the equation in finding the volume of a sphere. If I'm not mistaken I believe its (4/3) * pi * r^3 = Volume

OpenStudy (anonymous):

if you know the diameter the radius is half the diameter so r= 3 in

OpenStudy (anonymous):

i got 36pi

OpenStudy (anonymous):

Thats not it \[(4/3) \Pi 3^3\]

OpenStudy (anonymous):

\[(4/3) \Pi 9= 36/3 \Pi= 12\Pi\]

OpenStudy (anonymous):

thats the volume of the sphere but wait hold on. You want the volume of a hemisphere so its actually half a sphere. so the volume is \[6\Pi\]

OpenStudy (anonymous):

Tell me if you are still confused

OpenStudy (anonymous):

i got it thank and can u help me with other question

OpenStudy (anonymous):

Sure did you post it already?

OpenStudy (anonymous):

find the title that has a child's wading pool

OpenStudy (anonymous):

previous qstn: i got 18 pi

OpenStudy (anonymous):

volume for hemisfrera is (2/3)pi r^3 and r=3

OpenStudy (anonymous):

Actually Pika is right I ^2 instead of ^3 sorry we all make mistakes.

OpenStudy (anonymous):

that a okay